OpenStudy (anonymous):

[9.01] Which equation is a quadratic equation? y = −4(3x + 2) y + 3x3 = (x − 2)(3x + 8) y − 3 = (2x2 + 11)(x − 1) + 3x y + 5x = −2x(4 − x) + 1

OpenStudy (uri):

B,i think

OpenStudy (anonymous):

The second is quadratic equation :) y + 3x3 = (x − 2)(3x + 8) y+9=3x^2-8x-6x+16 y=3x^2-14x+16-9 y=3x^2-14x+7 Quadratic equation means x^2-x+6=0(This is an example of quadratic equation)